HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

LOW BARRIERS TO INTERNAL ROTATION IN THE MICROWAVE SPECTRUM OF 2,5-DIMETHYLFLUOROBENZENE

\underline{H. Y. SUN^{1}}, S. KHEMISSI^{2}, I. KLEINER^{1}, H. V. L. NGUYEN^{2,3},^{1}Université Paris Cité and Univ Paris Est Créteil, CNRS, LISA, Paris, France ,^{2}Univ Paris Est

Créteil and Université Paris Cité, CNRS, LISA, Créteil, France,^{3}Institut Universitaire de France (IUF), Paris, France

We investigated the rotational spectrum of 2,5-dimethylfluorobenzene which contains two coupled large amplitude motions of the methyl groups. The methyl internal rotations can be used to examine the molecular structure^{1} as steric and electronic environments around the methyl group affect its torsional barrier. Electronic properties play a particularly important role in aromatic molecules with the presence of a π-conjugated double bond system. 2,5-dimethylfluorobenzene was studied in the frequency range from 2 GHz to 26.5 GHz using a pulsed molecular jet Fourier transform microwave spectrometer. The experimental results are combined with quantumchemical calculations. The internal rotations of two inequivalent methyl groups with low torsional barriers (around 16 cm^{−1} and 226 cm^{−1}, respectively) lead to the splitting of all rotational transitions into quintets, resulting in separations of hundreds of MHz in the microwave spectrum. Spectral analysis and modeling of the observed splittings were performed using the programs XIAM^{2} and BELGI-C_{s}-2Tops,^{3} whereby the latter achieved measurement accuracy. The experimental results were compared with those of quantum chemistry calculations.

1doi:10.1002/cphc.201900400, M. Andresen, I. Kleiner, M. Schwell, W. Stahl, H. V. L. Nguyen, Sensing the Molecular Structures of Hexan-2-one by Internal Rotation and Microwave Spectroscopy, ChemPhysChem, 20, 2063-2073 (2019).

2doi:10.1515/zna-1996-0807, H. Hartwig and H. Dreizler, The Microwave Spectrum of trans-2,3-Dimethyloxirane in Torsional Excited States, Z. Naturforsch. 51a, 923-932 (1996).

3doi:10.1016/j.jms.2011.07.005, M. Tudorie, I. Kleiner, J.T.Hougen, S. Melandri, L.W. Sutikdja, W.

Stahl, A fitting program for molecules with two inequivalent methyl tops and a plane of symmetry at equilibrium: Application to new microwave and millimeter-wave measurements of methyl acetate, J.

Mol. Spectrosc. 269, 211-225 (2011).

p-number: p155  Submitted on Thu, 15 Jun 2023 10:53:43 +0200